OpenStudy (anonymous):

Simplify: 40-8(2x-1) I know it's very simple, but I'm not sure of my first step.

OpenStudy (mathstudent55):

Rewrite the 40. Then distribute the -8 by what is inside the parentheses. Then combine like terms.

OpenStudy (anonymous):

I'm not sure if I did it correctly, but my answer was 64x-32.

OpenStudy (mathstudent55):

Ok, here we go. One step at a time. 40 - 8(2x - 1) = 40 - 8 * 2x - 8(-1) = 40 - 16x + 8 = 48 - 16x

OpenStudy (mathstudent55):